b) two types of engines, A and B, were compared. Gas mileage, in miles per gallon, was measured. 50 experiments were conducted using engine type A and 75 experiments were done with engine type B. The gasoline used and other conditions were held constant. The average gas mileage was 36 miles per gallon for engine A and 42 miles per gallon for engine B. Find a 95% con dence interval on "\mu"B - "\mu"A, where "\mu"B and "\mu"Aare true mean gas mileages for engines A and B, respectively. population standard deviations are 6 and 8 for engines A and B, respectively. What can you conclude?


1
Expert's answer
2022-02-15T18:08:10-0500

Summary statistics:

Engine "A: \quad n_{1}=50 ; \quad \bar{x}_{1}=36 ; \quad \sigma_{1}=6"

Engine "B: \quad n_{2}=75 ; \quad \bar{x}_{2}=42 ; \quad \sigma_{2}=8"

Confidence level, "1-\alpha=95 \%=0.95"

Implies, level of significance, "\alpha=1-0.95=0.05"

Here, population standard deviations are known, so z-tabulated value "z_{\alpha / 2}" to find the confidence interval for "\mu_{1}-\mu_{2}" .

At "\alpha=0.05" , the two-tailed critical value from the z-table, "z_{\alpha / 2}=1.96"

Then, the 95% confidence interval for "\mu_{1}-\mu_{2}" is,

"\begin{aligned}

&\left(\bar{x}_{1}-\bar{x}_{2}\right) \pm z_{\alpha / 2} \sqrt{\frac{\sigma_{1}^{2}}{n_{1}}+\frac{\sigma_{2}^{2}}{n_{2}}} \\

&(36-42) \pm(1.96) \sqrt{\frac{6^{2}}{50}+\frac{8^{2}}{75}} \\

&-6 \pm 2.4585 \\

&(-8.4585,-3.5415)

\end{aligned}"

So, the 95% confidence interval on the difference of population mean gas mileages for engines A and B is "(-8.4585,-3.5415)."